# Sustainable Self-evolution Adversarial Training

### v1

@@ Abstract.
With the wide application of deep neural network models in various computer vision tasks, there has been a proliferation of adversarial example generation strategies aimed at deeply exploring model security. However, existing adversarial training defense models, which rely on single or limited types of attacks under a one-time learning process, struggle to adapt to the dynamic and evolving nature of attack methods. Therefore, to achieve defense performance improvements for models in long-term applications, we propose a novel Sustainable Self-Evolution Adversarial Training (SSEAT) framework. Specifically, we introduce a continual adversarial defense pipeline to realize learning from various kinds of adversarial examples across multiple stages. Additionally, to address the issue of model catastrophic forgetting caused by continual learning from ongoing novel attacks, we propose an adversarial data replay module to better select more diverse and key relearning data. Furthermore, we design a consistency regularization strategy to encourage current defense models to learn more from previously trained ones, guiding them to retain more past knowledge and maintain accuracy on clean samples. Extensive experiments have been conducted to verify the efficacy of the proposed SSEAT defense method, which demonstrates superior defense performance and classification accuracy compared to competitors.
@@
copyright: acmlicensedjournalyear: 2024doi: 10.1145/3664647.3681077conference: Make sure to enter the correct conference title from your rights confirmation email; October 28 - November 1, 2024; Melbourne, Australia.booktitle: Procedings ofthe 32nd ACM Intemational Conference on Mulimedia (MM’24), 0ctober 28-November 1, 2024, Melboume,VIC, Australiaisbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Computer vision tasksccs: Security and privacy Software and application security
@@ 1. Introduction
Deep learning has been widely applied in computer vision tasks such as image classification (Zhang et al., 2018; Chollet, 2017; Qian et al., 2019) and object detection (Long et al., 2023; Chen et al., 2023c; Qian et al., 2023), resulting in significant advancements. However, the vulnerability of deep learning models to adversarial attacks (Song et al., 2023; Kim et al., 2023a; Feng et al., 2023) has become a critical concern. Adversarial examples involve intentionally crafted small perturbations that deceive deep learning models, leading them to produce incorrect outputs. This poses a serious threat to the reliability and security of these models in real-world applications. For example, attackers can use adversarial glasses to mislead facial recognition systems. Consequently, research on defense mechanisms (Frosio and Kautz, 2023; Huang et al., 2023; Li and Xu, 2023) has become increasingly essential and urgent.For example, attackers can use adversarial glasses to mislead facial recognition systems.
@@
In this study, to address the above three challenges for the defense model against the ongoing generation of new adversarial examples, we propose a novel and task-driven Sustainable Self-Evolution Adversarial Training (SSEAT) framework to ensure the model maintains its accuracy and possesses robust and continuous defense capabilities. Our SSEAT defense method comprises three components: Continual Adversarial Defense (CAD), Adversarial Data Reply (ADR), and Consistency Regularization Strategy (CRS). To achieve sustainability in the adversarial training strategy (Challenge (1)), drawing inspiration from the continue learning paradigm, we propose a CAD pipeline, which learns from one type of adversarial example at each training stage to address the continual generation of new attack algorithms. To address the issue of catastrophic forgetting when continuously learning from various attacks (Challenge (2)), we introduce an ADR module to establish an effective re-learning sample selection scheme, advised by classification uncertainty and data augmentation. Meanwhile, to realize the trade-off between the model’s robustness against adversarial examples and accuracy on clean data (Challenge (3)), We design a CRS module to help the model not overfit to current attacks and prevent the model from losing knowledge on clean samples. Overall, our SSEAT method effectively addresses a range of defense challenges arising from continuously evolving attack strategies, maintaining high classification accuracy on clean samples, and ensuring lifelong defense performance against ongoing new attacks.
@@
  • We introduce a continual adversarial defense pipeline to learn from diverse types of adversarial examples across multiple stages, an adversarial data reply module to alleviate the catastrophic forgetting problem when the model continuously learns from new attacks, and a consistency regularization strategy to prevent significant accuracy drop on clean data samples.
@@ 2.2. Adversarial Training
Figure 2. (a) Illustration of our Continual Adversarial Defense (CAD) pipeline. CAD helps the model to learn from new kinds of attacks in multiple stages continuously. (b) Illustration of our Adversarial Data Replay (ADR) module. ADR guides the model in selecting diverse and representative replay data to alleviate the catastrophic forgetting issue. (c) Illustration of our Consistency Regularization Strategy (CRS) component. CRS encourages the model to learn more from the historically trained models to maintain classification accuracy.
@@ 2.3. Contiune Learning
Continue learning (Kirkpatrick et al., 2017; Ayub and Wagner, 2023; Wei et al., 2023) aims at the model being able to continuously learn new data without forgetting past knowledge. Continue learning can be divided into three main categories. One is based on the regularization of model parameters (Kim et al., 2023b; Zaken et al., 2021) by preserving important parameters from the past while updating less critical ones. However, this method’s performance is not ideal when applied to scenarios involving a large number of tasks. One is based on knowledge playback (Lin et al., 2023; Nie et al., 2023; Chrysakis and Moens, 2020; Sun et al., 2022; Bang et al., 2021), where important past samples are stored in memory and used for training when encountering new tasks. Another approach involves dynamically expanding model parameters (Hung et al., 2019; Li et al., 2019; Ge et al., 2023) to assign different parameters to different tasks. This method helps alleviate catastrophic forgetting and enhances model performance, but it requires significant memory and computational resources. Nevertheless, continue learning models are also susceptible to adversarial examples. Wang et. al. (Wang et al., 2023a) attempt to combine adversarial training with continue learning paradigms, but they do not consider the defensive performance of the current model against old attacks in long-term application scenarios. Differently from previous work, our focus is on ensuring that when the model continues to encounter new adversarial attacks, it can maintain robustness against past adversarial examples and improve resilience against new adversarial attacks.
@@ 3. Methodology
3.1. Task Definition and Framework Overview
@@ 3.2. Continual Adversarial Defense
[…]
@@ 3.3. Adversarial Data Reply
where $\mathcal{D}_{t}$ represents the sample set of the $t_{th}$ learning stage in CAD task.
@@
[…]
@@
We assess the sample’s uncertainty to the perturbation by,
@@
[…]
where $r\left(x\right)$ represents the uncertainty of sample $x$, $Q_{c}$ indicates the number of times augmented samples are predicted as the true class, $\mathcal{W}$ represents the one-hot encoded class vector, where only the element corresponding to the true class is true. Lower values of $r\left(x\right)$ indicate that the sample resides closer to distribution center.
We allocate memory for a replay buffer of size $K$ for each learning iteration. We sort all samples $\mathcal{D}_{t}$ based on the computed uncertainty $r\left(x\right)$, and sample the examples with an interval of $\left|\mathcal{D}_{t}\right|/K$. We have diversified the replay samples by sampling perturbed samples of varying intensities, ranging from robust to fragile ones, which can broaden the scope of memories, encompassing a wide range of complex scenarios.
@@ 3.4. Consistency Regularization Strategy
Considering data points ($\left(x,y\right)$ drawn from distribution $\mathcal{D}$, and augmentations $A_{1}$ and $A_{2}$ sampled from set $\mathcal{A}$ , we denote the adversarial noise of $A_{i}\left(x\right)$ as $\delta_{i}$ It is obtained by $\delta_{i}:=argmax_{\left\|\delta\right\|_{p\leq\epsilon}}\\ \mathcal{L}\left(A_{i}\left(x\right),y,\delta;\theta_{t}\right)$. Our objective is to regularize the temperature-scaled distribution $\hat{f_{\theta_{t}}}\left(x;\tau\right)$ of adversarial examples across augmentations for consistency. Here, $\tau$ is the temperature hyperparameter. Specifically, we use temperature scaling to adjust the classifier: $\hat{f_{\theta_{t}}}\left(x;\tau\right)=Softmax\left(\frac{z_{\theta_{t}}\left(x\right)}{\tau}\right)$, where $z_{\theta_{t}}\left(x\right)$ is the logit value of $f_{\theta_{t}}\left(x\right)$ before the softmax operation. Therefore, we obtain the regularization loss as follows:
@@ 4.1. Experimental Settings
Implementation Details. We summarize the training procedure of our SSEAT framework in Alg. 1. The model uses resnet18 (He et al., 2016) as the classification network structure. We implement Torchattacks (Kim, 2020) to generate 100 adversarial images per category for each adversarial attack strategy, for a total of 1000 images. On the CIFAR-10 and CIFAR-100 dataset , before training, each image is resized to 32×32 pixels, and underwent data augmentation, which included horizontal flipping and random cropping. For the training hyperparameters of experiments, we use a batch size of 8, the epoch for training clean samples is set to 40, and the epoch for training adversarial samples is set to 20. We train the model using the SGD optimizer with momentum 0.9 and weight decay $5\times 10^{-4}$. In addition, we set the memory buffer size to 1000.
Competitors. We compare our SSEAT method with several adversarial training works, such as the PGD-AT (Madry et al., 2017), TRADES(Zhang et al., 2019), MART(Wang et al., 2019), AWP (Wu et al., 2020), RNA (Dong et al., 2022), LBGAT (Cui et al., 2021) and RIFT(Zhu and G., 2023).
@@ 4.2. Experimental Results
We have conducted extensive experiments over the CIFAR-10 and CIFAR-100 datasets with various attack orders in the CDS task, compared to several competitors and baseline. All attack results are reported under the black-box condition.
For the CDS task, our SSEAT model can achieve the best robustness against ongoing generated new adversarial examples. To better evaluate our SSEAT model under the CDS task, We report model robustness over various attack orders compared to several competitors over two datasets. As shown in Tab. 1, Tab. 2, and Tab. 3, our SSEAT method can beat all adversarial training competitors for all adversarial examples, which demonstrate the efficacy of SSEAR framework to tackle the continuous new attacks under black-box condition.
For the CDS task, our SSEAT model can maintain competitive classification accuracy over the clean data. For real-life applications, in addition to continuously improving the defense performance against new attack methods, the model also needs to have good recognition effects on clean samples. Thus, we report the classification accuracy over original data in Tab. 1, Tab. 2, and Tab. 3. The results show our SSEAT method not only achieves strong robustness under complex changes, but also has good performance on the original task. our SSEAT successfully defends against different attacks while maintaining good performance on clean samples. It clearly reveals the advantage of SSEAT in terms of generalization.Achieving a balanced trade-off between robustness and accuracy greatly enhances the practicality of the adversarial training strategy in real scenarios.
For the CDS task, our SSEAT can realize the best defense results on unknown attacks. Without loss of generality, we use the SSEAT model, trained on CIFAR-10. Interestingly, our SSEAT outperforms other competitors on all unknown attack defenses as in Tab. 4. It strongly points out the generalized robustness of our SSEAT method.
Table 4. Comparing results against unknown adversarial attacks under CDS task.
@@
Table 5. Results obtained from several variants of our SSEAT.
Method FGSM BIM PGD RFGSM NIM Clean
Baseline (CAD) 65.71 72.58 71.80 72.84 71.16 63.30
+CRS 70.78 73.97 73.89 73.20 71.96 80.22
+ADR 72.58 75.81 75.60 75.23 74.31 81.97
+ADR+CRS (Ours) 74.47 76.53 76.97 77.66 76.13 82.92
4.3. Ablation Study
The efficacy of each component in the proposed method. As shown in Tab. 5, by comparing the results of different variants, we notice the following observations, (1) ‘Baseline (CAD)’ based on the basic continue learning pipeline, can only realize learning from continuous attacks, however, the performance on clean samples is too poor to meet the actual application scenarios; (2) Comparing the results between ‘+Baseline (CAD)’ and ‘+CRS’, obviously, we can notice the classification accuracy over the clean samples can be obviously improved. CRS can better handle the trade-off between model robustness and accuracy; (3) Comparing ‘+Baseline (CAD)’ and ‘ADR’, such results illustrate that our sampling strategy is effective, and can select representative and diverse samples at each training stage; (4) Comparing the ‘+CRS’, ‘+ADR’, and ‘+ADR+CRS (Ours)’, the results clearly demonstrate the ADR and CRS modules can complement each other on both classification accuracy and defense robustness.
perturbation FGSM PGD20 PGD50 SIM DIM Clean
8/255 71.6/74.9 78.0/78.4 75.8/76.1 60.6/75.6 70.4/72.3 83.0/83.1
@@ 4.3. Ablation Study
Table 8. The results of our SSEAT with different numbers of training data in one training stage in the CAD pipeline. The first column “K” in the table represents the amount of training data used in adversarial training at each stage.
@@
Figure 3. Visualization of clean examples representations for CAD baseline and SSEAT by using t-SNE(Linderman et al., 2017). We use 1000 test images from the CFAIR-10 dataset for visualization. Different colors represent different categories.

### v2

@@ Abstract.
With the wide application of deep neural network models in various computer vision tasks, there has been a proliferation of adversarial example generation strategies aimed at exploring model security deeply. However, existing adversarial training defense models, which rely on single or limited types of attacks under a one-time learning process, struggle to adapt to the dynamic and evolving nature of attack methods. Therefore, to achieve defense performance improvements for models in long-term applications, we propose a novel Sustainable Self-evolution Adversarial Training (SSEAT) framework. Specifically, we introduce a continual adversarial defense pipeline to realize learning from various kinds of adversarial examples across multiple stages. Additionally, to address the issue of model catastrophic forgetting caused by continual learning from ongoing novel attacks, we propose an adversarial data replay module to better select more diverse and key relearning data. Furthermore, we design a consistency regularization strategy to encourage current defense models to learn more from previously trained ones, guiding them to retain more past knowledge and maintain accuracy on clean samples. Extensive experiments have been conducted to verify the efficacy of the proposed SSEAT defense method, which demonstrates superior defense performance and classification accuracy compared to competitors.
@@
copyright: acmlicensedjournalyear: 2024doi: https://doi.org/10.1145/3664647.3681077conference: Make sure to enter the correct conference title from your rights confirmation email; October 28 - November 1, 2024; Melbourne, Australia.isbn: 978-1-4503-XXXX-X/18/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
@@ 1. Introduction
Deep learning has been widely applied in computer vision tasks such as image classification (Zhang et al., 2018; Chollet, 2017; Qian et al., 2019) and object detection (Long et al., 2023; Chen et al., 2023c; Qian et al., 2023), resulting in significant advancements. However, the vulnerability of deep learning models to adversarial attacks (Song et al., 2023; Kim et al., 2023a; Feng et al., 2023) has become a critical concern. Adversarial examples involve intentionally crafted small perturbations that deceive deep learning models, leading them to produce incorrect outputs. This poses a serious threat to the reliability and security of these models in real-world applications. Consequently, research on defense mechanisms (Frosio and Kautz, 2023; Huang et al., 2023; Li and Xu, 2023) has become increasingly essential and urgent.
@@
In this study, to address the above three challenges for the defense model against the ongoing generation of new adversarial examples, we propose a novel and task-driven Sustainable Self-evolution Adversarial Training (SSEAT) framework to ensure the model maintains its accuracy and possesses robust and continuous defense capabilities. Our SSEAT defense method comprises three components: Continual Adversarial Defense (CAD), Adversarial Data Reply (ADR), and Consistency Regularization Strategy (CRS). To achieve sustainability in the adversarial training strategy (Challenge (1)), drawing inspiration from the continue learning paradigm, we propose a CAD pipeline, which learns from one type of adversarial example at each training stage to address the continual generation of new attack algorithms. To address the issue of catastrophic forgetting when continuously learning from various attacks (Challenge (2)), we introduce an ADR module to establish an effective re-learning sample selection scheme, advised by classification uncertainty and data augmentation. Meanwhile, to realize the trade-off between the model’s robustness against adversarial examples and accuracy on clean data (Challenge (3)), We designed a CRS module to help the model not overfit to current attacks and prevent the model from losing knowledge on clean samples. Overall, our SSEAT method effectively addresses a range of defense challenges arising from continuously evolving attack strategies, maintaining high classification accuracy on clean samples, and ensuring lifelong defense performance against ongoing new attacks.
@@
  • We introduce a continual adversarial defense pipeline to learn from diverse types of adversarial examples across multiple stages, an adversarial data reply module to alleviate the catastrophic forgetting problem when the model continuously learns from new attacks, and a consistency regularization strategy to prevent significant accuracy drop on clean data.
@@ 2.2. Adversarial Training
Figure 2. (a) Illustration of our Continual Adversarial Defense (CAD) pipeline. CAD helps the model to learn from new kinds of attacks in multiple stages continuously. (b) Illustration of our Adversarial Data Replay (ADR) module. ADR guides the model to select diverse and representative replay data to alleviate the catastrophic forgetting issue. (c) Illustration of our Consistency Regularization Strategy (CRS) component. CRS encourages the model to learn more from the historically trained models to maintain classification accuracy.
@@ 2.3. Contiune Learning
Continue learning (Kirkpatrick et al., 2017; Ayub and Wagner, 2023; Wei et al., 2023) aims at the model being able to continuously learn new data without forgetting past knowledge. Continue learning can be mainly divided into three categories. One is based on the regularization of model parameters (Kim et al., 2023b; Zaken et al., 2021) by preserving important parameters from the past while updating less critical ones. However, this method’s performance is not ideal when applied to scenarios involving a large number of tasks. One is based on knowledge playback (Lin et al., 2023; Nie et al., 2023; Chrysakis and Moens, 2020; Sun et al., 2022; Bang et al., 2021), where important past samples are stored in memory and used for training when encountering new tasks. Another approach involves dynamically expanding model parameters (Hung et al., 2019; Li et al., 2019; Ge et al., 2023) to assign different parameters to different tasks. This method helps alleviate catastrophic forgetting and enhances model performance, but it requires significant memory and computational resources. Nevertheless, continue learning models are also susceptible to adversarial examples. Wang et. al. (Wang et al., 2023a) attempt to combine adversarial training with continue learning paradigms, but they do not consider the defensive performance of the current model against old attacks in long-term application scenarios. Differently from previous work, our focus is on ensuring that when the model continues to encounter new adversarial attacks, it can maintain robustness against past adversarial examples and improve resilience against new adversarial attacks.
@@ 3. Methodology
3.1. Task Defination and Framework Overview
@@ 3.2. Continual Adversarial Defense
In a classification task, a dataset $\mathcal{D}$ consists of $n$ pairs $(x_{i},y_{i})$, where $x_{i}$ represents input samples and $y_{i}$ denotes corresponding class labels ranging from integers $1$ to $K$. The classification model $f_{\theta}$ is intended to map the input space $\mathcal{X}$ to the output space $\Delta^{K-1}$, generating probability outputs through a softmax layer. To deal with the boom-growing attack strategies, the concept of adversarial robustness extends beyond evaluating the model’s performance solely on $\mathcal{P}$. It involves assessing the model’s ability to handle perturbed samples within a certain distance metric range around $\mathcal{P}$. Specifically, our goal is to achieve $\mathit{l_{p}-robustness}$, where we aim to train a classifier $f_{\theta}$ to accurately classify samples $\left(x+\delta,y\right)$ under any $\delta$ perturbation such that $\left\|\delta\right\|_{p}\leq\epsilon$. Here, $\left(x,y\right)$ follows the distribution $\mathcal{P}$, and $p\geq 1$ with a small $\epsilon>0$.
@@ 3.3. Adversarial Data Reply
where $\mathcal{D}_{t}$ represents the sample set of the $t_{th}$ learning stage in CAD.
@@
$p\left(y=c|x\right)=\int_{\tilde{\mathcal{D}}}p\left(y=c|\tilde{x}_{t}\right)p\left(\tilde{x}_{t}|x\right)d\tilde{x}_{t}$
@@
We assess the sample’s uncertainty in relation to the perturbation by,
@@
(5) $r\left(x\right)=1-\frac{1}{A}\underset{c}{max}Q_{c}$
where $r\left(x\right)$ represents the uncertainty of sample $x$, $Q_{c}$ indicates the number of times augmented samples are predicted as the true class, $\mathcal{W}$ represents the one-hot encoded class vector, where only the element corresponding to the true class is true. Lower values of $r\left(x\right)$ indicate that the sample resides closer to the distribution center.
We allocate memory for a replay buffer of size $K$ for each learning iteration. We sort all samples $\mathcal{D}_{t}$ based on the computed uncertainty $r\left(x\right)$, and sample the examples with an interval of $\left|\mathcal{D}_{t}\right|/K$. We have diversified the replay samples by sampling perturbed samples of varying intensities, ranging from robust to fragile ones, which can broaden the scope of memories, encompassing a wide range of scenarios.
@@ 3.4. Consistency Regularization Strategy
Considering data points ($\left(x,y\right)$ drawn from distribution $\mathcal{D}$, and augmentations $A_{1}$ and $A_{2}$ sampled from set $\mathcal{A}$ , we denote the adversarial noise of $A_{i}\left(x\right)$ as $\delta_{i}$ . It is obtained by $\delta_{i}:=argmax_{\left\|\delta\right\|_{p\leq\epsilon}}\\ \mathcal{L}\left(A_{i}\left(x\right),y,\delta;\theta_{t}\right)$. Our objective is to regularize the temperature-scaled distribution $\hat{f_{\theta_{t}}}\left(x;\tau\right)$ of adversarial examples across augmentations for consistency. Here, $\tau$ is the temperature hyperparameter. Specifically, we use temperature scaling to adjust the classifier: $\hat{f_{\theta_{t}}}\left(x;\tau\right)=Softmax\left(\frac{z_{\theta_{t}}\left(x\right)}{\tau}\right)$, where $z_{\theta_{t}}\left(x\right)$ is the logit value of $f_{\theta_{t}}\left(x\right)$ before the softmax operation. Therefore, we obtain the regularization loss as follows:
@@ 4.1. Experimental Settings
Implementation Details. We summarize the training procedure of our SSEAT framework in Alg. 1. The model uses resnet18 (He et al., 2016) as the classification network structure. We implement Torchattacks (Kim, 2020) to generate 100 adversarial images per category for each adversarial attack strategy, for a total of 1000 images. On the CIFAR10 dataset, before training, each image is resized to 32×32 pixels, and underwent data augmentation, which included horizontal flipping and random cropping. For the training hyperparameters of experiments, we use a batch size of 8, the epoch for training clean samples is set to 40, and the epoch for training adversarial samples is set to 20. We train the model using the SGD optimizer with momentum 0.9 and weight decay $5\times 10^{-4}$. In addition, we set the memory buffer size to 1000.
Competitors. We compare our SSEAT method with several adversarial training works, such as the PGD-AT (Madry et al., 2017), TRADES(Zhang et al., 2019), MART(Wang et al., 2019), AWP (Wu et al., 2020), RNA (Dong et al., 2022), and LBGAT (Cui et al., 2021),RIFT(Zhu and G., 2023).
@@ 4.2. Experimental Results
We have conducted extensive experiments over the CIFAR-10 dataset with various attack orders in the CDS task, compared to several competitors and baseline. All attack results are reported under the black-box condition.
For the CDS task, our SSEAT model can achieve the best robustness against ongoing generated new adversarial examples. To better evaluate our SSEAT model under the CDS task, We report model robustness over various attack orders compared to several competitors. As shown in Tab. 1, Tab. 3, and Tab. LABEL:cifar100_1, our SSEAT method can beat all adversarial training competitors for all adversarial examples, which demonstrate the efficacy of SSEAR framework to tackle the continuous new attacks under black-box condition.
For the CDS task, our SSEAT model can maintain competitive classification accuracy over the clean data. For real-life applications, in addition to continuously improving the defense performance against new attack methods, the model also needs to have good recognition effects on clean samples. Thus, we report the classification accuracy over original data in Tab. 1, Tab. 3, and Tab. LABEL:cifar100_1. The results show our SSEAT method not only achieves strong robustness under complex changes, but also has good performance on the original task. Achieving a balanced trade-off between robustness and accuracy greatly enhances the practicality of the adversarial training strategy in real scenarios.
Results on unkown attacks. Without loss of generality, we use the SSEAT model, trained on CIFAR-10. Interestingly, our SSEAT outperforms other competitors on all unknown attack defenses as in Tab. 6. It strongly points out the generalized robustness of our SSEAT method.
Table 4. Results obtained from several variants of our SSEAT model.
Method FGSM BIM PGD RFGSM NIM Clean
Baseline (CAD) 65.71 72.58 71.80 72.84 71.16 63.30
+CRS 70.78 73.97 73.89 73.20 71.96 80.22
+ADR 72.58 75.81 75.60 75.23 74.31 81.97
+ADR+CRS (Ours) 74.47 76.53 76.97 77.66 76.13 82.92
perturbation FGSM PGD20 PGD50 SIM DIM Clean
8/255 71.6/74.9 78.0/78.4 75.8/76.1 60.6/75.6 70.4/72.3 83.0/83.1
Table 6. Defense results against unknown adversarial attacks
@@
Table 8. The results of our SEAT with different numbers of training data in one training stage in the CAD pipeline. The first column “K” in the table represents the amount of training data used in adversarial training at each stage.
@@
Figure 3. Visualization of clean examples representations for CAD and SSEAT by using t-SNE(Linderman et al., 2017). We use 1000 test images from CFAIR-10 dataset for visualization. Different colors represent different categories.
4.3. Ablation Study
The efficacy of each component in the proposed method. As shown in Tab. 4, by comparing the results of different variants, we notice the following observations, (1) ‘Baseline (CAD)’ based on the basic continue learning pipeline, can only realize learning from continuous attacks, however, the performance on clean samples is too poor to meet the actual application scenarios; (2) Comparing the results between ‘+Baseline (CAD)’ and ‘+CRS’, obviously, we can notice the classification accuracy over the clean samples can be obviously improved. CRS can better handle the trade-off between model robustness and accuracy; (4) Comparing ‘+Baseline (CAD)’ and ‘ADR’, such results illustrate that our sampling strategy is effective, and can select representative and diverse samples at each training stage; (4) Comparing the ‘+CRS’, ‘+ADR’, and ‘+ADR+CRS (Ours)’, the results demonstrate the ADR and CRS modules can complement each other on both classification accuracy and defense robustness;
To better understand the insight of our proposed model and verify its effectiveness, we further conduct several ablation studies below to evaluate different design choices.
Visualization experiments of clean samples representations. We additionally visualize the representations of clean examples with CAD and SSEAT. As shown in Fig. 3, we can see that compared with CAD, SSEAT exhibits more distinct distances between classes in feature representations on clean samples, with closer intra-class distances, which enhances its discriminative ability.For example, in SSEAT, the distribution of red points representing Class 0 is more clustered, whereas in CAD, the distribution is more dispersed. This observation indicates that SSEAT obtains more discriminative representations of clean examples compared with CAD. These results indicate that our SSEAT can maintain high accuracy on clean samples.